# Graph Network Simulator MGN MeshGraphNet

### v1

@@ 3.1 Denoising Diffusion Probalistic Models (DDPMs) for time-dependent simulations
As discussed in Section 3.1, early denoising prediction steps resemble one-step models, while later steps progressively refine spatial frequency bands, from coarse structures to fine details. After $m<K$ denoising steps at time $t$, the intermediate prediction $\tilde{\mathbf{u}}^{t}_{0|m}$ thus already captures low-to-mid frequency information. We assume this information is sufficient to condition the next step $t{+}1$, partiularly to predict a solution within this already refined frequency band. Using these properties, we begin denoising of time step $t{+}1$ after $m$ steps by initializing $\mathbf{u}^{t{+}1}_{K}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, allowing batched denoising of steps $t{:}t{+}1$. More generally, after $hm$ denoising steps, we initialize the time step $t+h$. We denoise a rolling prediction window of $t:t+h$ time steps in parallel. After warm up, this process yields a constant prediction window of size $h=K/m$ and reduces the number of model calls from $KT$ to $K-m+mT$ steps. We denote $m$ as denoising stride.
Figure 2 b) visualizes this for the special case $m=1$ and $K=3$ diffusion steps. Each model call moves the simulation forward in time by one step, denoising each of the $K$ partially denoised predictions in its rolling windows by one additional noise step. A denoising step of ROBI consists of two consecutive steps. First the learned model predicts the denoising velocity $\mathbf{v}_{\theta}(\mathbf{u}^{t_{\text{B}}}_{k_{\text{B}}},k_{\text{B}},\mathbf{u}^{t_{\text{B}}-1}_{0})$ of the prediction horizon batch with the time steps $t_{\text{B}}=t:t+h$ and denoising steps $k_{\text{B}}=m:m:K$. Subsequently, the denoising scheduler is used to reconstruct the clean samples $\tilde{\mathbf{u}}^{t_{b}}_{0|k_{b}}$ and to update the noisy samples $\mathbf{u}^{t_{\text{B}}}_{k_{\text{B}}-1}\sim p_{\theta}(\mathbf{u}^{t_{\text{B}}}_{k_{\text{B}}-1}|\mathbf{u}^{t_{\text{B}}}_{k_{\text{B}}})$. After each model call with a batch size of $h=3$, the prediction window shifts one physical time step forward.
@@
For $m=K$ROBI converges to conventional autoregressive diffusion inference (cf. Figure 2 a)). Thus, the denoising stride $m$ can be considered a hyperparameter that trades off prediction accuracy and memory usage with inference speed. Notably, ROBI does not affect the training process or the model architecture. The model remains a one-step diffusion model, which preserves the time-shift equivariance of Markovian systems, fast training convergence and small GPU memory utilization. Furthermore, we only require the last time step as history, which proves to be more stable and accurate for autoregressive ML-based simulations [5, 9].
In practice, we find that denoising residual predictions $\Delta\mathbf{u}^{t}_{k}$ instead of $\mathbf{u}^{t}_{k}$ improves accuracy. Let $\Delta\tilde{\mathbf{u}}^{t_{\text{B}}}_{0}\approx\Delta\tilde{\mathbf{u}}^{t_{\text{B}}}_{0|k_{\text{B}}}$ be the predicted clean residuals with $t_{\text{B}}=t:t+h$ and denote the last fully denoised state as $\mathbf{u}^{t-1}_{0}$. We extend ROBI to update the clean state estimate $\tilde{\mathbf{u}}^{t+j}_{0}$ of the time step $t+j$, $j\in\{0,h\}$ in the prediction window using the cumulative sum $\tilde{\mathbf{u}}^{t+j}_{0}=\mathbf{u}^{t-1}_{0}+\sum_{i=0}^{j}\Delta\tilde{\mathbf{u}}^{t+i}_{0}$.
@@
3.2 Denoising Diffusion Probalistic Models (DDPMs) for mesh-based simulations
@@
To fully utilize Denoising Diffusion Probalistic Models (DDPMs)’s potential for generating rich, multi-frequency solutions, prediction models must handle multi-scale information. Hierarchical Graph Neural Networks (HGNNs) are particularly well-suited for this, as their architecture inherently learns representations at varying levels of granularity, mirroring the diverse frequency content present in DDPM outputs. Leveraging this idea, we train HGNN to predict the discrete denoising velocity $\mathbf{v}_{i,\theta}(\mathbf{u}^{t}_{i,k},k,\mathbf{u}^{t-1}_{i,0})$ for mesh-based simulations on the mesh graph $\mathcal{G}=(\mathcal{V},\mathcal{E}^{\text{M}})$. It takes the current noisy sample $\mathbf{u}^{t}_{i,k}$ and is conditioned on the previous clean sample $\mathbf{u}^{t-1}_{i,0}$.
Root-node AMG-based Mesh Coarsening. To apply HGNN, we first create a hierarchical mesh graph $\mathcal{G}^{0:L}=(\mathcal{V}^{0:L},\mathcal{E}^{0:L,\text{M}})$ consisting of $L+1$ graphs with nodes $\mathcal{V}^{0:L}$ and mesh edges $\mathcal{E}^{0:L,\text{M}}$ using the mesh graph $\mathcal{G}$. We define the fine mesh graph $\mathcal{G}^{0}:=\mathcal{G}$ as our mesh graph at level $l=0$ and coarsen it $L$ times. For coarsening, we use an algebraic root-node-based smooth aggregation solver [53] implemented in [59]. Starting with the adjacency matrix (with self-loops) of the fine mesh $A^{0}$, the solver creates a hierarchy of adjacency matrices $A^{0:L}$.
@@
We use root-node-based smoothed aggregation, which leverages algebraic information from the adjacency matrix, to construct coarse graphs $\mathcal{G}^{1:L}$ and inter-level transfer operators. The selected root nodes $j\in\mathcal{V}^{1:L}$, a subset of the fine mesh nodes, form the coarse graph nodes. We reuse the smoothed aggregation mapping from the AMG algorithm, assigning each fine node $i\in\mathcal{V}^{l}$ to exactly one root node $j\in\mathcal{V}^{l+1}$, to define coarse graph edges. Following [52], we add an edge between $i,j\in\mathcal{V}^{l+1}$ if their aggregation nodes $\mathcal{V}^{l}$ of graph $\mathcal{G}^{l-1}$ are connected. Unlike bi-stride pooling with Delaunay remeshing [22], this AMG-based approach better preserves mesh geometry by leveraging strength connection in the adjacency matrix. Figure 3 visualizes this difference.
@@
We additionally define $L-1$ downsampling edges $\mathcal{E}^{l,\text{D}}$ and upsampling edges $\mathcal{E}^{l,\text{U}}$, which connect the nodes of the graphs $\mathcal{G}^{l}$ and $\mathcal{G}^{l+1}$, and vice versa. We define these as the connections (non-zero values) of the restriction and prolongation matrices of the AMG hierarchy, resulting in an extended hierarchy graph $\mathcal{G}^{0:L}=(\mathcal{V}^{0:L},\mathcal{E}^{0:L,\text{M}}\cup\mathcal{E}^{0:L-1,\text{D}}\cup\mathcal{E}^{0:L-1,\text{U}})$. Those matrices are constructed by smoothing the sparse initial aggregation mapping using the adjecency matrix, resulting in a denser matrix with a wider receptive field. As before, we found that the obtained pooling mappings respect the mesh geometry well, as shown in Figure 3 d).
Figure 3: Comparison of AMG-based mesh coarsening to BSDL. a) the original, fine mesh. b) the mesh after two BSDL coarsening steps. c) the mesh after one AMG coarsening steps. This mesh has approximately as many nodes as the mesh in b). d) up- and downsampling edges from level 00 to $1$ (top) and level $1$ to $2$. Bright blue indicates coarse nodes.
@@
For contact experiments, we further extend the graph hierarchy $\mathcal{G}^{0:L}$ by adding contact edges $\mathcal{E}^{0,\text{C}}$ [22] between nodes. In a simulation involving two colliding components, we define a bidirectional edge between the nodes of the first part and the nodes of the second part if their distance is less than the specified contact radius $R$. With these edges, we obtain a hierarchy $\mathcal{G}^{0:L}=(\mathcal{V}^{0:L},\mathcal{E}^{0:L,\text{M}}\cup\mathcal{E}^{0:L,\text{C}}\cup\mathcal{E}^{0:L-1,\text{D}}\cup\mathcal{E}^{0:L-1,\text{U}})$.
@@
Encoder. Let $\mathcal{G}^{0:L}$ be a hierarchical graph as defined above and $\mathbf{u}^{t}_{i,k}$ the noisy sample at denoising step $k$ of simulation step $t$. We define node embeddings $\mathbf{k}_{i}\in\mathcal{V}^{0:L}$, mesh edge embedding $\mathbf{e}_{ij}^{\text{M}}\in\mathcal{E}^{0:L,\text{M}}$, contact edge embeddings $\mathbf{e}_{ij}^{\text{C}}\in\mathcal{E}^{\text{C}}$, downsampling edge embeddings $\mathbf{e}^{\text{D}}_{ij}\in\mathcal{E}^{0:L-1,\text{D}}$ and upsampling edge embeddings $\mathbf{e}^{\text{U}}_{ij}\in\mathcal{E}^{0:L-1,\text{U}}$. We add relative node distances $\mathbf{x}^{t}_{ij}=\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}$ and their norm $|\mathbf{x}^{t}_{ij}|$ to all edge embeddings and the initial distance $\mathbf{x}^{0}_{ij}=\mathbf{x}^{0}_{i}-\mathbf{x}^{0}_{j}$ and their norm $|\mathbf{x}^{0}_{ij}|$ to mesh edges, down- and upsampling embeddings. Node embeddings include a one-hot encoding $n_{i}$ of the node type. Node embeddings at level $l=0$ additionally include $\mathbf{u}^{t}_{i,k}$ and task-specific features. All embeddings are projected to the latent dimension $d$ via linear layers. We add a Fourier encoding [57] for the denoising step $k$ and the normalized level $l^{*}=l/L$ to inform the AMPN of relative graph depth.
Processor and Decoder. Similar to AMG solvers [54, 53] and UNets [60]Algebraic-hierarchical Message Passing Networks (AMPNs) use a V-cycle to propagate information between levels. They consist of five core message passing modules: Pre-Processing, Downsampling, Solving, Upsampling and Post-Processing, as shown in Figure 1. Pre-Processing, Solving and Post-Processing modules use an  Intra-Level-Message Passing Stack (Intra-MP-Stack) consisting of $N$ message passing steps to update the heterogeneous subgraph $\mathcal{G}^{l}=(\mathcal{V}^{l},\mathcal{E}^{l,\text{C}}\cup\mathcal{E}^{l,\text{M}})$ of level $l$. Given node embeddings $\mathbf{k}_{i}\in\mathcal{V}^{l}$, contact edge embeddings $\mathbf{e}_{ij}^{\text{C}}\in\mathcal{E}^{l,\text{C}}$ and mesh edge embeddings $\mathbf{e}_{ij}^{\text{M}}\in\mathcal{E}^{l,\text{M}}$, the message passing update of the level graph at step $n$ is defined by
$\mathbf{e}_{ij}^{\text{C},n+1}$ $=W^{n}_{\theta,\mathcal{E}^{\text{C}}}~{}\mathbf{e}_{ij}^{\text{C},n}+f^{n}_{\theta,\mathcal{E}^{\text{C}}}(\mathbf{k}^{n}_{i},\mathbf{k}^{n}_{j},\mathbf{e}_{ij}^{\text{C},n})\textrm{ , }$ (2)
$\mathbf{e}_{ij}^{\text{M},n+1}$ $=W^{n}_{\theta,\mathcal{E}^{\text{M}}}~{}\mathbf{e}_{ij}^{\text{C},n}+f^{n}_{\theta,\mathcal{E}^{\text{M}}}(\mathbf{k}^{n}_{i},\mathbf{k}^{n}_{j},\mathbf{e}_{ij}^{\text{M},n})\textrm{ , }$
@@
$\mathbf{e}_{ij}^{n+1}$ $=W^{n}_{\theta,\mathcal{E}}~{}\mathbf{e}_{ij}^{n}+f^{n}_{\theta,\mathcal{E}}(\mathbf{k}^{\text{rec},n}_{i},\mathbf{k}^{\text{send},n}_{j},\mathbf{e}_{ij}^{n})\textrm{ , }$ (3)

### v2

@@ 3.1 Denoising Diffusion Probabilistic Models (DDPMs) for time-dependent simulations
As discussed in Section˜3.1, early denoising steps behave similarly to one-step model predictions, while later steps progressively refine higher spatial frequencies (from coarse structures to fine details). After $m<K$ denoising steps at time $t$, the intermediate estimate $\tilde{\mathbf{u}}^{t}_{0|m}$ already captures low-to-mid frequency content that is sufficient to condition the next physical step t+1t{+}1 to predict a solution within this already refined frequency band.
ROBI exploits this property by starting the denoising of step t+1t{+}1 as soon as step $t$ has progressed by $m$ steps. We initialize $\mathbf{u}^{t{+}1}_{K}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and denoise the batch of both time steps $t_{w}{\in}\{t,t+1\}$. More generally, after $jm$ denoising steps, we initialize the time step t+jt{+}j and denoise a rolling time window with $t_{w}{\in}\{t,...,t+j\}$ in parallel. After a short warm-up, fully denoised samples (those with k=0k{=}0) drop out and the window size becomes constant with w=K/mw{=}K/m and $t_{w}{\in}\{t,...,t{+}w{-}1\}$. Notably, the denoising index $k$ and thus the noise level increase along the window toward later physical times, aligning with the natural growth of forecast uncertainty. This reduces the number of model calls from $KT$ (conventional inference) to Km+mTK{-}m{+}mT steps. For m=1m{=}1, the number of model calls reduces to K+TK{+}T, which is effectively the same complexity $\mathcal{O}(T)$ as for one-step models, when $K\ll T$. We refer to $m$ as the denoising stride. Figure˜2 b) visualizes the special case m=1m{=}1 and K=3K{=}3 diffusion steps. Each model call advances the simulation by one physical step and applies one denoising step to each of the $K$ partially denoised predictions in the window.
A single Batched Denoising Step (cf. Figure˜2 b) left for m=1m{=}1 and K=3K{=}3) of ROBI consists of two consecutive steps. In the Prediction step, the model outputs the velocity $\mathbf{v}_{\theta}(\mathbf{u}^{t_{w}}_{k_{w}},k_{w},\mathbf{u}^{t_{w}-1}_{0})$ for each element in the prediction window $t_{w}{\in}\{t,...,t{+}w{-}1\}$, with the diffusion indices $k_{w}={\in}\{m,...,K\}$. Subsequently, we perform a State Update step. Using the scheduler, we reconstruct $\tilde{\mathbf{u}}^{t_{w}}_{0|k_{w}}$ to update the conditioning states $\mathbf{u}^{t_{w}}_{0}$, and sample the reverse transition $\mathbf{u}^{t_{w}}_{k_{w}-1}\sim p_{\theta}(\mathbf{u}^{t_{w}}_{k_{w}-1}|\mathbf{u}^{t_{w}}_{k_{w}})$. Both are used in the next Prediction Step as inputs. After $m$ such Batched Denoising Steps, we advance in time and drop the fully denoised states (k=0k{=}0) and initialize a new Gaussian sample (k=Kk{=}K) for the next time index t+wt{+}w, so the subsequent call evaluates $\mathbf{v}_{\theta}(\mathbf{u}^{t_{w}+1}_{k_{w}},k_{w},\mathbf{u}^{t_{w}}_{0})$.
@@
For m=Km{=}KROBI reduces to conventional autoregressive diffusion inference (cf. Figure˜2 a)). Thus, the denoising stride $m$ can be considered a hyperparameter that trades off prediction accuracy and memory usage with inference speed. Most notably, in State Update, ROBI reconstructs the physical states, while the rolling time window is treated purely as a batch dimension in the prediction model. Consequently, the model always predicts $\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k)=\mathbf{v}_{\theta}(\mathbf{u}^{t}_{k},k,\mathbf{u}^{t-1}_{0})$, i.e., conditioned only on the previous reconstructed state during both training and inference, which proves to be more stable and accurate for autoregressive ML-based simulations [5, 9]. Furthermore, this preserves the time-shift equivariance of Markovian systems, fast training convergence and small GPU memory utilization.
In practice, we find predicting residuals $\Delta\mathbf{u}^{t}_{k}$ improves accuracy. Let the model denoise a batch of residual states $\Delta\tilde{\mathbf{u}}^{t_{w}}_{0}\approx\Delta\tilde{\mathbf{u}}^{t_{w}}_{0|k_{w}}$ of the time window $t_{w}{\in}\{t,...,t{+}w{-}1\}$ and denote the last fully denoised state as $\mathbf{u}^{t-1}_{0}$. We then recover clean states inside the window via a cumulative sum $\tilde{\mathbf{u}}^{t+j}_{0}=\mathbf{u}^{t-1}_{0}+\sum_{i=0}^{j}\Delta\tilde{\mathbf{u}}^{t+i}_{0},~j\in\{0,...,w-1\}$. To further accelerate inference, we optionally stop denoising early at a truncation step $k_{\mathrm{tr}}$ and use the partially denoised state $\tilde{\mathbf{u}}^{t}_{0|k_{\text{tr}}}$ as the final prediction of the time step. This remains effective because early denoising steps already approximate one-step prediction.
@@
3.2 Denoising Diffusion Probabilistic Models (DDPMs) for mesh-based simulations
@@
To fully utilize Denoising Diffusion Probabilistic Models (DDPMs)’s potential for generating rich, multi-frequency solutions, prediction models must handle multi-scale information. Hierarchical Graph Neural Networks (HGNNs) are particularly well-suited for this, as their architecture inherently learns representations at varying levels of granularity, mirroring the diverse frequency content present in DDPM outputs. Leveraging this idea, we train HGNN to predict the discrete denoising velocity $\mathbf{v}_{i,\theta}(\mathbf{u}^{t}_{i,k},k,\mathbf{u}^{t-1}_{i,0})$ for mesh-based simulations on the mesh graph $\mathcal{G}=(\mathcal{V},\mathcal{E}^{\text{M}})$. It takes the current noisy sample $\mathbf{u}^{t}_{i,k}$ and is conditioned on the previous clean sample $\mathbf{u}^{t-1}_{i,0}$.
@@
Root-node AMG-based Mesh Coarsening. We construct a hierarchical mesh graph $\mathcal{G}^{\text{H}}=\mathcal{G}^{0:L}=(\mathcal{V}^{0:L},\mathcal{E}^{0:L,\text{M}})$ consisting of $L+1$ mesh graphs with nodes $\mathcal{V}^{0:L}$ and mesh edges $\mathcal{E}^{0:L,\text{M}}$ by coarsening the fine graph $\mathcal{G}^{0}:=\mathcal{G}$ $L$ times. Coarsening is performed with root-node smoothed aggregation [53] implemented in [59]. The solver takes the fine-mesh adjacency (with self-loops) $A^{0}$ as its system matrix and creates a hierarchy of adjacency $A^{0:L}$, upsampling (prolongation) $U^{0:L-1}$ and downsampling (restriction) $D^{0:L-1}$ matrices. In practice, our implementation only requires the fine adjacency $A^{0}$, the root-node solver [59], and constructing the graphs of each level from the non-zeros of the returned matrices $A^{1:L}$. The resulting nodes $j\in\mathcal{V}^{0:L}$ of each level are always a subset of the fine mesh nodes. Unlike bi-stride pooling with Delaunay remeshing [22], this AMG-based approach better preserves mesh geometry by leveraging the strength of connections in the adjacency matrix. Figure˜10 in Appendix˜E visualizes this difference.
@@
We additionally define L1L-1 downsampling edges $\mathcal{E}^{l,\text{D}}$ and upsampling edges $\mathcal{E}^{l,\text{U}}$, which connect nodes between successive levels $\mathcal{G}^{l}$ and $\mathcal{G}^{l+1}$. We define these as the connections (non-zero values) of the given up- $U^{0:L-1}$ and downsampling $D^{0:L-1}$ matrices, resulting in an extended hierarchy graph $\mathcal{G}^{\text{H}}=(\mathcal{V}^{0:L},\mathcal{E}^{0:L,\text{M}}\cup\mathcal{E}^{0:L-1,\text{D}}\cup\mathcal{E}^{0:L-1,\text{U}})$. As before, we found that the obtained pooling mappings respect the mesh geometry well, as shown in Figure˜10 d) in Appendix˜E.
@@
For contact experiments, we add contact edges $\mathcal{E}^{0,\text{C}}$ [22] to the graph hierarchy $\mathcal{G}^{\text{H}}$. In a simulation involving two colliding components, we define a bidirectional edge between the nodes of the first part and the nodes of the second part if their distance is less than the specified contact radius $R$. The resulting hierarchy is given by $\mathcal{G}^{\text{H}}=(\mathcal{V}^{0:L},\mathcal{E}^{0:L,\text{M}}\cup\mathcal{E}^{0:L,\text{C}}\cup\mathcal{E}^{0:L-1,\text{D}}\cup\mathcal{E}^{0:L-1,\text{U}})$.
@@
Encoder. Let $\mathcal{G}^{\text{H}}$ be a hierarchical graph as defined above and $\mathbf{u}^{t}_{i,k}$ the noisy sample at denoising step $k$ of simulation step $t$. We define node embeddings $\mathbf{k}_{i}\in\mathcal{V}^{0:L}$, mesh edge embedding $\mathbf{e}_{ij}^{\text{M}}\in\mathcal{E}^{0:L,\text{M}}$, contact edge embeddings $\mathbf{e}_{ij}^{\text{C}}\in\mathcal{E}^{\text{C}}$, downsampling edge embeddings $\mathbf{e}^{\text{D}}_{ij}\in\mathcal{E}^{0:L-1,\text{D}}$ and upsampling edge embeddings $\mathbf{e}^{\text{U}}_{ij}\in\mathcal{E}^{0:L-1,\text{U}}$. We add relative node distances $\mathbf{x}^{t}_{ij}=\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}$ and their norm $|\mathbf{x}^{t}_{ij}|$ to all edge embeddings and the initial distance $\mathbf{x}^{0}_{ij}=\mathbf{x}^{0}_{i}-\mathbf{x}^{0}_{j}$ and their norm $|\mathbf{x}^{0}_{ij}|$ to mesh edges, down- and upsampling embeddings. Node embeddings include a one-hot encoding $n_{i}$ of the node type. Node embeddings at level $l=0$ additionally include $\mathbf{u}^{t}_{i,k}$ and task-specific features. All embeddings are projected to the latent dimension $d$ via linear layers. We add a Fourier encoding [57] for the denoising step $k$ and the normalized level $l^{*}=l/L$ to inform the AMPN of relative graph depth.
Processor and Decoder. Similar to AMG solvers [54, 53] and UNets [60]Algebraic-hierarchical Message Passing Networks (AMPNs) use a V-cycle to propagate information between levels. They consist of five core message passing modules: Pre-Processing, Downsampling, Solving, Upsampling and Post-Processing, as shown in Figure˜1. Pre-Processing, Solving and Post-Processing modules use an  Intra-Level-Message Passing Stack (Intra-MP-Stack) consisting of $N$ message passing steps to update the heterogeneous subgraph $\mathcal{G}^{l}=(\mathcal{V}^{l},\mathcal{E}^{l,\text{C}}\cup\mathcal{E}^{l,\text{M}})$ of level $l$. Given node embeddings $\mathbf{k}_{i}\in\mathcal{V}^{l}$, contact edge embeddings $\mathbf{e}_{ij}^{\text{C}}\in\mathcal{E}^{l,\text{C}}$ and mesh edge embeddings $\mathbf{e}_{ij}^{\text{M}}\in\mathcal{E}^{l,\text{M}}$, the message passing update of the level graph at step $n$ is defined by
[…]
@@
[…]